OpenStudy (anonymous):

6x+4=5-6x been a while since I did this, need a bit of help..

OpenStudy (anonymous):

I know the answer is 1/12, I just can't remember the correct steps to get there

OpenStudy (anonymous):

move the common variables to one side

OpenStudy (anonymous):

so, 6x+6x= 5-4

OpenStudy (anonymous):

12x=1

OpenStudy (anonymous):

wait, so 12x=1?

OpenStudy (anonymous):

yep, then isolate the x

OpenStudy (anonymous):

divide the 12 on both sides

OpenStudy (anonymous):

OKay, so then it would be x=1/12 o-o that was so simple, thanks!

OpenStudy (anonymous):

no prob!